evaluate the following integral
∫ 1 / x log x dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello student,
Please find the answer to your question below

I = \int \frac{1}{xlogx}dx
logx = t
\frac{dx}{x} = dt
I = \int \frac{1}{t}dt
I = log(t) + c
I = log(logx) + c